OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (anonymous):

divide by -4 first

OpenStudy (anonymous):

so u get |x + 5| = 4

OpenStudy (anonymous):

from this since there an absolute there are 2 x answer 4 and -4

OpenStudy (anonymous):

u solve the rest

OpenStudy (anonymous):

x+5 = 4 or x+5 = -4

OpenStudy (anonymous):

So it would be x = −1, x = 9

OpenStudy (anonymous):

What the heck? X = -1, x = 9

OpenStudy (mathmath333):